A map has a scale of 2 inches to every 25 miles. How far does 6 inches on the map represent in miles? Use a proportion to solve.
Answer by info@email-tutors.com(9)

Hello!
Proportions are in fraction form, so here goes:
2 inches / 25 miles = 6 inches / x miles
cross multiply to get: (2 inches )* (x miles ) = (6 inches) * (25 miles)
You will get: 2 * x = 6 * 25
You will then get: 2 * x = 150
divide both sides by two to get: 75 as the value of "x"
This means that 6 inches is equivalent to 75 miles
For GUARANTEED math homework help for just $1.00 (one dollar) per problem (with a swift response), check out our website (view profile) or go to email hyphen tutors.com